Andy Denis City University, London

Holism versus reductionism in economic thought A response to Mario Bunge

ABSTRACT.This paper is a response to Mario Bunge (2000) ŚSystemism: the alternative to individualism and holismŠ. By compressing the two overlaying oppositions, between holism and reductionism and between idealism and materialism, into a single spectrum running from ŚindividualismŠ via ŚsystemismŠ to ŚholismŠ, Bunge loses focus and clarity on the fundamental division between holistic and reductionistic world-views in economics and social science in general. This lack of clarity itself contributes to an over-simplification and misunderstanding of the relation between methodological premises and policy consequences. Alongside the reductionist free marketeers of the kind Bunge criticises, such as Friedman, Lucas and Sir Keith Joseph, there are also holist <i>laissezfaire</i> theorists such as Smith and Hayek. Ignoring the methodological differences between them is both mistaken and allows us to overlook the fundamentally ideological role of invisible hand mechanisms in allowing economists to retain some approximation to efficiency as their default notion of how the economy actually works.

Keywords: Holism; reductionism; systems thinking; idealism; materialism; economics; policy prescription; invisible hand.

1. Introduction

This paper is a response to Mario Bunge (2000) ŚSystemism: the alternative to individualism and holismŠ. Bunge’s paper is an extremely interesting article which makes a number of telling points and is evidence of a growing discomfort with the reductionism of the neoclassical school currently hegemonic within the there are two major points which need to made with respect to the article.Firstly, it is, in my opinion, preferable explicitly to distinguish between the two oppositions to which Bunge is implicitly referring: reductionism versus holism, on the one hand, and materialism versus idealism on the other. The second point is that the relation between policy prescription and philosophical standpoint in political economy is both more complex and more fascinating than BungeŠs account would lead one to believe.

Section 2 of the paper considers the first of these points and proceeds to make some initial comments on the opposition between reductionism and holism in economics. Section 3 looks in more detail at the second question, that of the relationship between methodological standpoint and policy prescription in political economy. Section 4 concludes.

2. Holism versus reductionism and materialism versus idealism

This section of the paper distinguishes the two oppositions with which, I suggest, BungeŠs article is really concerned, and goes on to discuss the significance of one of these, that between reductionism and holism, in political economy.

Bunge’s thesis is that there are three fundamental research approaches in the social sciences: the two most influential, individualism and holism, being fatally flawed, with only the minority approach of systemism offering a viable way forward. The first two are inadequate Ű each managing to avoid the otherŠs error only by committing an opposite error of its own, while systemism manages to synthesise the other two, accepting the criticism each makes of the other. For Bunge, ŚholismŠ is inadequate Śbecause there are no relations without relataŠ; and ŚindividualismŠ so Śbecause all individuals are interrelatedŠ(147)

Neither of the two most influential approaches to the study and management of social affairs is completely adequate. Individualism is deficient because it underrates or even overlooks the bonds among people, and holism, because it plays down or even enslaves individual action. By contrast, systemism makes room for both agency and structure.Ť (156-157)

Systemism, apart from being defined negatively with respect to individualism and holism, is characterised by viewing the world from a systems perspective: Śeverything is either a system or a component of a system and every system has peculiar (emergent) properties that its components lackŠ (147).

Now, there is nothing in these formulations which is incorrect and the points made are valuable ones. However, I think greater clarity can be obtained by stating the matter slightly differently. We are fully justified in endorsing BungeŠs rejection of ŚholismŠ as ignoring the fact that the relations, of which our systems of relations consist, depend fundamentally upon the properties of the <i>relata</i>, the substrate-level entities which are doing the relating.ă But in doing so we take a position on a fundamental opposition: that between materialism and idealism. When we observe, and investigate, some facet of the world, we can choose Ű assuming we are at least monists, of one kind or another Ű to be led by one of two kinds of expectation: that the phenomenon in question is fundamentally reducible to matter in motion, or, on the contrary, that the properties of the entity in question derive ultimately from the mind, whether the mind of the individual observer, or that of some more general consciousness such as a deity. In rejecting what he calls ŚholismŠ, Bunge is taking his stand with materialism, and opposing idealism. What Bunge describes as ŚholismŠ makes the properties of an entity independent of the properties of its material substrate.

We are equally justified in complaining, with Bunge, that ŚindividualismŠ fails to take into consideration the <i>relationships</i> between agents, the fact that individuals are only nodes in <i>systems</i> of such relationships, and the <i>emergence</i> properties at the macro, or system level. But when we do so, we are again taking a position on a particular opposition which has implicitly or explicitly underlain a vast amount of methodological discourse in economics and elsewhere: that between holism and reductionism. We can see things in the world, and, in particular, economic phenomena, as congeries of substrate-level entities Ű or we can attempt to understand them as a whole or system, with emergent properties not enjoyed by constituent components.

So, on the materialism-idealism polarity, we may see that BungeŠs systemism adopts a materialist standpoint, while on the clash between holism and reductionism, he supports holism A better Ű clumsier, perhaps, but more exact title for BungeŠs article would therefore have been ŚThe systems standpoint or materialist holism: the alternative to reductionism and idealismŠ.

Of these two oppositions, I am most immediately concerned with that between holism and reductionism. Materialism has underpinned the natural scientific revolutions of the past half-millennium, and thereby gained enormous prestige (thank goodness). That, on the one hand, it has done so in spite of a considerable admixture of reductionism, and, on the other, holism has so often been associated with idealism (the contributions of Hegel, Toynbee and Smuts being prime examples Ű see, for example, Hegel, 1975, Toynbee, 1972, and Smuts, 1996), have both contributed to the peculiarly low esteem, indeed, contempt, in which holism is held today. The list of systems thinkers Ű scientists, philosophers, and evolutionary theorists working within a materialistic holism Ű who have felt it necessary to condemn ŚholismŠ is lengthy indeed. ă Dawkins (1989: 113), Minsky (1985: 27), and Dennett (1995: 80-82) are just a few of the most salient examples Ű and, of course, Bunge himself. So, while idealism is largely (though certainly not entirely) discredited, reductionism is in far better health. I want, therefore, to say more on this second opposition, and will start with two proposed definitions:

Reductionism – the view that an entity at one level can be understood as a congeries, an aggregate of entities at a lower, substrate level, that the properties and behaviour of higher level entities can be understood in terms of the properties and behaviour of its constituent lower level parts, <i>taken in isolation</i>.

Holism – the view that phenomena at one level can be understood as emergent at that level, that a higher level entity can be understood as a product of the <i>interrelationships</i> between its component parts.

Correspondence relating to previous versions of the paper suggest that some discussion is in order at this point. Firstly, the question has been raised (Bunge, personal communication) as to whether ŚreductionismŠ and holismŠ so defined are ontological or epistemological standpoints As stated, both definitions are epistemological Ű Śan entity Ě can be understood as ĚŤ. Nevertheless, if we adopt a realist standpoint, as in fact I do, then the implication is that a phenomenon can be understood in a particular way precisely because that is the way it is. ă So as well as being explicitly epistemological, the definitions are implicitly ontological. However, that further step is unnecessary for the purposes of this paper. ă

Secondly, the exact definitions of the terms are as controversial as everything else in the debate between the supporters of each point of view. The above definitions are my own, and are as unlikely to please as any others Correspondents, responding to earlier versions of this paper, have criticised this or that statement about reductionism or holism in the paper but because their criticism was on the basis of other definitions than those set out above, they failed to engage with the points that I am making submit that what is critical here is that holism and reductionism so defined can be shown to characterise two living trends in economic thought and to throw up interesting and enlightening questions about the nature of those trends. ă

Keynes, for example, in the PrefaceŠ to the French edition of 1939 of the <i>General Theory</i>, clearly advocates a holistic approach: I have called my theory a <i style=’mso-bidi-font-style: normal’>general </i>theory. ă I mean by this that I am chiefly concerned with the behaviour of the economic system as a whole .... And I argue that important mistakes have been made through extending to the system as a whole conclusions which have been correctly arrived at in respect of a part of it taken in isolation.Ť (Keynes, 1973: xxxii)

Keynes sets out very clearly here what he takes to be the distinguishing feature of the two approaches: that, on the one hand, we can derive <b>correct</b> conclusions from the study of <i>micro</i>economic phenomena Śtaken in isolationŠ, but that to extend those conclusions to <i style=’mso-bidi-font-style: normal’>macro</i>economic phenomena leads to error, and, on the other, that the correct approach is (what we would now call) a systems approach, aiming to examine the behaviour of Śthe economic system as a wholeŠ. The fallacy of composition involved in the approach he is criticising here, derives from a difference, between the two levels, in what it is legitimate to take as parametric or given (Keynes, 1973: 293).</p>

Robert Lucas, on the contrary, is a very clear spokesman for the trend in economics which favours a reductionist methodology. ă The following is taken from the final paragraph of his <i>Models of Business Cycles</i> (Lucas, 1987: 107): The most interesting recent developments in macroeconomic theory seem to me describable as the reincorporation of aggregative problems such as inflation and the business cycle within the general framework of ŚmicroeconomicŠ theory. ă If these developments succeed, the term Śmacro-economicŠ will simply disappear from use and the modifier ŚmicroŠ will become superfluous. We will simply speak, as did Smith, Ricardo, Marshall and Walras, of economic </i>theory. If we are honest, we will have to face the fact that at any given time there will be phenomena that are well-understood from the point of view of the economic theory we have, and other phenomena that are not. ă We will be tempted, I am sure, to relieve the discomfort induced by discrepancies between theory and facts by saying that the ill-understood facts are the province of some other, different kind of economic theory. ă Keynesian ŚmacroeconomicsŠ was, I think, a surrender (under great duress) to this temptation. ă It led to the abandonment Ě of the use of the only Śengine for the discovery of truthŠ that we have in economics.

Here we have a clear expression of the desire to reduce macroeconomics to microeconomics, and a characterisation of the Keynesian approach as an illegitimate Śsurrender to temptationŠ.

Both the passages cited occur in contexts Ű a preface, and the concluding paragraph of a book Ű where the authors are standing back from the detail of the theories that they are presenting, and indicating what they regard as the underlying general features of their approaches. ă What they choose to highlight in both cases is their selection of a holist or reductionist approach. ă This, I think, establishes, at least a prima facie </i>case, that the issue is worth looking at and potentially useful in understanding the controversy between various schools of thought in the history of economics. ă

Some examples can be given to show how this controversy continually emerges in economics: in an extended review of Binmore (1994), comments thus on the latter’s deployment of evolutionary game theory Ű and it should be born in mind that what he says is intended to be, not praise, but a very serious criticism of BinmoreŠs approach: with [BinmoreŠs] new definition of self-interest, methodological individualism, a hallmark of neoclassical economics, is abandoned. ă In its place is a more holistic and integrated view of society as an integral organismŤ.

Demonstrating the dependence of all macroeconomics on microeconomic principles is essential for the fulfillment of the (methodological) individualist requirements of neoclassical economicsŤ (cited in Nelson, 1984: 576) So Dore and Boland agree that the reductionist standpoint Ű Śmethodological individualismŠ Ű is an essential component of neoclassical economics. ă In a similar vein, Schotter (1985: 2), right at the beginning of a book on <i>Free Market Economics</i>, says that in the libertarian, or classical liberal view, Śsociety is nothing more than the aggregate of the individuals composing itŠ. ă That is, before anything else, Schotter feels it important to establish the reductionist standpoint of libertarianism.

Hoover (1995), on the other hand, in a special edition of <i>The Monist</i> on The Metaphysics of Economics, argues that The idea that macroeconomics stands in need of a microfoundational base is a commonplace among economists Ě [but] ontological reduction of macroeconomics to microeconomics is untenable Ě while the program of microfoundations may illuminate macroeconomics in various ways, it cannot succeed in its goal of replacing macroeconomics.Ť

Feeling it necessary to point out that an example she wants to use for purposes of illustration, although couched in reductionist terms, could be replaced by a more holistic version, Nancy Cartwright remarks that We could attempt to explain macroeconomic regularities by reference to capacities and relations that can only sensibly be attributed to institutions or to the economy as a whole, with no promise of reduction to features of individuals.Ť ă (1995: 278)

Nelson (1984) is an interesting case since at first he seems enthusiastic about the possibility of reducing macro to micro in economics: The possibility of reducing at least parts of macroeconomic theory to microeconomic theory is especially exciting Ě. [I]t seems that the largeŰscale phenomena dealt with in macroeconomics must be the results of the total effects of the small scale phenomena dealt with in microeconomics. ă Therefore, one might expect that bridges could be built by merely adding up the microeconomic laws describing the microphenomena to obtain the macroeconomic laws describing macrophenomena Ě. [I]t does seem that a reduction of macroeconomics to microeconomics would not be plagued with the kind of ontological difficulties which might attend [other reduction] schemesŤ (Nelson, 1984: 573-74).

His conclusion, however, is pessimistic. ă There is little prospect of providing anything which deserves to be called microfoundations Ě it is unlikely that we will see any progress in reducing macroeconomics to microeconomics.Ť (ibid: 593)

The way forward Ě is to eschew the reductionist search for finer-grain microfoundations and instead [to] study how complex economic systems can emerge from the interactions between agentsŤ (Cross and Strachan, 1997: 565). Finally, Hayek, who spent much of his life criticising the linked errors of macroeconomics, economic statistics and socialism, was ă amongst the most trenchant of those wishing to deny the legitimacy of macroeconomics: ŚI recognize Ě microeconomic theory as the only legitimate economic theory (1983: 22); ŚI believe it is only microeconomics which enables us to understand the crucial functions of the market processŠ (Hayek, 1983: 27); Ś[M]y disagreement with that book [sc Keynes’s <i>General Theory</i>] did not refer so much to any detail of the analysis as the general approach followed in the whole work. ă The real issue was the validity of ... macro-analysisŠ. ă(Hayek, 1978: 100). ă As we shall see, however, the citation of these passages by no means settles the question of Hayek’s stance on the methodological issue of reductionism versus holism. ăă

This section has argued that BungeŠs triad individualism-systemism-holism rests on a pair of more fundamental dyads Ű reductionism-holism and idealism-materialism.ă Further, I have proposed definitions of reduction and holism and tried to show that the opposition between them is a living issue in economics. ă The next section takes up the issue of the association, which can be seen to emerge here, and which is asserted by Bunge, between methodological premises and policy prescription.

3.Policy prescription and social philosophy

A pattern seems to emerge from the examples cited: apparently there is a tendency for libertarians and those on the right of the spectrum of policy views within economics, such as Lucas, to appeal to reductionist methodological premises, while those on the left, those like Keynes adopting a more interventionist stance, are more likely to invoke holistic underpinnings for their theoretical pronouncements. Bunge takes up this point, speaking of the obsolete individualism of Ě Smith Ě [and] the neoclassical economistsŠ (147); the cultural policy of liberalism, which is based on individualism, is one of benign neglect. ă By contrast, the totalitarian cultural policy, which is based on holism, is one of censorshipŠ (150-151); Śall market worshippers espouse individualismŠ (151).ŚThe radical individualists oppose all social planning in the name of individual liberties Ě holists swear by top-down planning Ě they are likely to ignore their aspirations and rights [sc those of the common people]. ă In either case, the powerless individual, whether forsaken or corralled, has nothing to gainŠ (153). ă By contrast, Śsystemism takes into account social values (ignored by individualism) as well as individual values (ignored by holism)Š (157). ă

But there is a paradox here. ă We have just seen that there <i>seems </i>to be an association between a <i>laissez-faire</i> policy prescription and a reductionist methodology. ă Nevertheless, starting with Adam Smith a profoundly influential trend, epitomised by writers such as Friedrich Hayek and Armen Alchian, has proposed what has been called an Śinvisible hand theorem (De Vany, 1996: 427), in which order in human affairs is ŚemergentŠ, Śresulting from human action but not designŠ. ă This emergence seems flatly to contradict the association between laissez-faire and reductionism just noted. ă This section will explore that paradox. ă

Perhaps the canonical statement of the Śinvisible hand theoremŠ is SmithŠs statement that It is not from the benevolence of the butcher, the brewer, or the baker, that we expect our dinner, but from their regard to their own interest. ă We address ourselves, not to their humanity but to their self-love, and never talk to them of our own necessities, but of their advantages.Ť ă

Even here, in this well-known and apparently simple statement, there is something mysterious about the relation between micro and macro levels. ă The butcher, brewer and baker do not care about the dinners they provide, so that, in some sense, the desirable social outcome of feeding the members of society is achieved in spite of rather than because of the motives and behaviours of the food providers. ă The articulation between the motivation and behaviour of agents at the micro level and macro outcomes is not even at first blush a trivial or straightforward question. It seems highly paradoxical that the trend which is associated with this Śinvisible hand theoremŠ, a trend which I describe as ŚreductionistŠ, should refer to the ŚemergentŠ properties of human order, when emergence is precisely what distinguishes the holist from the reductionist approach.

The question therefore arises, whether the invisible hand theorem is consistent with the reductionist or the holist approach. ă Haworth (1994 Ch 4, Section 1 ŚA logical dilemmaŠ, 32-35) performs an invaluable service by addressing precisely this issue. ă HaworthŠs procedure Ű part of his philosophical critique of libertarian thinking Ű is as follows. ă Firstly, he identifies two ŚthesesŠ, implicit in libertarian thought, which he defines as follows, each illustrated by a statement from a libertarian source. ă Then the libertariansŠ logical dilemma arises from the mutual incompatibility of the two theses:

<b>The reducibility thesis</b>: the fully developed market economy can be understood as the sum or aggregate of its discrete components, the individual bilateral exchanges at the micro level. ă Sir Keith Joseph: ŚSince inequality arises from the operation of innumerable preferences, it cannot be evil unless those preferences are themselves evil.

<b>The invisible hand thesis</b>: the market is a Śparadigmatic exemplarŠ of want-satisfaction (ie, unrestricted market forces leave agents better off than any alternative economic environment) because an invisible hand transmutes our self-interested behaviour into socially desirable outcomes. Mandeville: Śthe grand principle that makes us social creatures, the solid basis, the life and support of all trade and employment without exception is

So reductionism says that evil only comes of evil, while the job of the invisible hand is specifically to transmute evil into good. For Sir Keith the aggregate outcome cannot be evil as long as the preferences it is based on are innocent; for Mandeville, on the contrary, the aggregate outcome cannot be good unless the preferences underpinning it are evil, vicious, selfish. ă Thus Haworth is able to conclude that the libertarians cannot have it both ways: Ślibertarianism is seriously broken backed in the sense that it must abandon one of its central theses.Š (Haworth, 1994: 34)

My own approach is slightly different from HaworthŠs. ă Firstly, we need to pay more attention to the association, alluded to earlier, between methodological standpoint and policy prescription. There are policy implications of the choice between reductionism and holism. ă And, indeed, the consequences for policy implied by the approach selected, so far from being a mere scholium, are the tail which wags the methodological dog.

If one adopts the systems approach and recognises that the unintended collective outcomes of an unplanned, uncoordinated mass of individual actions may have far from desirable features, then the obvious implication is to see whether there is anything we can do about it. ă The absence of an invisible hand invites the intervention of the very visible hand of state intervention. The reductionist approach, on the contrary, says that, assuming individuals can be counted on to do the best they can for themselves given the constraints they face, the aggregate outcome of those individual actions will also be the best available: state intervention in the economy is nugatory.

There are two possibilities: we could be living in a world where the reductionists are right or one where the holists are right. ă Needless to say, I think we reside in the latter. ă If we lived in the former, the macro level would simply reflect the micro level.ă There would be nothing for an invisible hand to do. ă The individual would be directly social, or, what comes to the same thing, there would be no separate category of the social. ă Individual utility maximisation would directly be social welfare maximisation: the distinction between them would be meaningless. ă Likewise, macro irrationality would be just a summary of micro irrationality: unemployment would either be a product of irrational behaviour by workers, such as Śpricing themselves out of jobsŠ, or it would be the product of a rational desire for leisure and, hence, itself rational. In general, individuals could with confidence be left to get on with it without supervision or intervention. ă A reductionist world would be a laissez-faire world.

If, on the other hand, we were to inhabit, as in my opinion we do, a holistic world then reductionists would (and do) have a problem. It is fairly obvious that higher level entities are not simply aggregates of their micro components: water does not behave as an aggregate of hydrogen and oxygen; steam, liquid water, and ice do not consist of tiny gaseous, liquid and solid molecules; nor do chairs consist of hard, green, ugly or uncomfortable molecules (Haworth, 1994: 35). As Ken Binmore (1994: 213) says, ŚAs we all know, you would be very lucky if a small change in your program led only to a small change on your screen. ă Tiny programming errors typically lead to wild and unpredictable results when the program is run.ŠAll these properties emerge at higher levels. ă The problem faced by the reductionist is how to reconcile this fact of an obvious disjuncture between levels with the laissez-faire policy prescription. ă Libertarians face severe difficulties sustaining a logically consistent reductionism in a holistic world. ă

The invisible hand is one potential solution to this problem. ă There are two possibilities. ă Either one can ignore the disjuncture between levels, and adopt a thoroughgoing reductionist methodology and policy stance Ű this seems to be line taken by Joseph, Lucas and Friedman Ű or with Hayek and Adam Smith one can accept that disjuncture and adopt a methodological holism but at the same time postulate a mechanism reconciling that methodological holism with a <i>laissez-faire</i> policy reductionism. ă Such a mechanism is an invisible hand mechanism. The invisible hand allows us to say, granted that social outcomes are not logically bound to reflect individual behaviour in an aggregative, summary manner, nevertheless a mechanism exists which ensures that in practice </i>they do so. ă The invisible hand is what allows us to think, and act, in a reductionist way in a holistic world: it underpins reductionism by tacitly conceding holism. Laissez-faire</i> is vindicated, we are inveigled into tying the visible hand behind our back, if we can be persuaded that the invisible hand will do its job instead, and do it better.

What I am suggesting, therefore, is the following: the <i>laissez-faire</i> policy prescription does, indeed, embody a reductionist standpoint. ă However, there is more than one way of sustaining that standpoint methodologically. ă One can believe, or at least act as if one believes, that the world truly is reductionist in relevant ways and that supposed macro-level pathology is simply the summation of micro-level behaviour which may or may not be pathological. <i>Laissez-faire</i> is a reductionist policy prescription in the sense that it issues from a reductionist methodological standpoint. ă Or one can accept that the world is holistic and hence that macro-level pathologies might in principle be emergent at that level, but postulate the existence of an invisible hand mechanism which ensures that the reductionist policy prescription of <i>laissez-faire</i> is nevertheless valid. The latter strategy combines methodological holism with policy reductionism.

Bunge’s criticisms of the ŚindividualismŠ (for which, following the terminology I adopt, read <i>reductionism</i>) of the neoclassical economists are well taken - although it may well be a rhetorical error blandly to dismiss monetarist macroeconomics as akin to faith healing (156)! ă But the simple, one-to-one relationship between this reductionist standpoint and a <i>laissez-faire</i> policy prescription, which Bunge asserts, simply doesnŠt exist. Compare the standpoints of Hayek and Friedman. ă For Friedman, economics is based on the study of Śa number of independent households – a collection of Robinson CrusoesŠ (cited in Haworth, 1994: 8). For Hayek, on the contrary, Śindividuals are merely the foci in the network of relationshipsŠ (Hayek, 1979: 59). So on the definitions proposed earlier, Hayek subscribes to a very clearly holistic, and Friedman to an equally clearly reductionist methodology. Yet they still both endorse the same basic framework for policy prescription: <i>laissez-faire</i>. ă And in bracketing Smith with the neoclassicals (148) in the reductionist camp, Bunge is simply in error Ű see Denis (1999, 2000).

Smith and Hayek are methodologically very distant from the crude reductionism of Joseph, Lucas and Friedman. They tacitly recognise a holistic world by invoking invisible hand mechanisms. ă For Smith the invisible hand is literally the hand of an omniscient and omnipotent deity desiring nothing other than the maximisation of human welfare. For Hayek, writing in a more secular age, the invisible hand mechanism takes the form of an evolutionary process, albeit, unfortunately, one based on the exploded group selection theory of VC Wynne-Edwards. And HayekŠs attempts to distinguish his own stance from that of Keynes, his anxiety to head off a line of thought leading from the holistic or systems thinking premises he shared with Keynes to an interventionist policy prescription, lead him to make the crudely reductionistic statements about macroeconomics and microeconomics which we noted earlier.

The resolution, then, to the paradox identified at the beginning of this section is as follows. We can understand the contradictory association between an appeal to reductionist methodological underpinnings and an assertion of a holistic invisible hand mechanism on the basis of a dual foundation. ă The first part is a factual hypothesis about the nature of the world we actually inhabit, namely, the systems view of the world, the hypothesis that the world is a holistic one with all that this implies about the scope and potential for collective intervention in the economy. ă The second part concerns the laissez-faire policy prescription of the reductionist camp, a policy prescription which both depends on and supports the reductionist methodological stance. Taking the two points together, we can see that some mechanism has to be introduced to articulate between a holist world and a reductionist policy prescription: the invisible hand does the job. The invisible hand, whatever its precise content, whether it comprises the hand of a deity as in Smith, or the result of an evolutionary process as in Hayek, allows us to assert that, in practice, the world can be treated as if it were reductionist. The logical inconsistency which Howarth has correctly identified is an internalisation of the inconsistency between programme and holistic world. ă It is illogical because it is attempting to do what is ultimately impossible, namely to reconcile the irreconcilable. ă

The alternative to both of these approaches is to combine recognition of the holistic nature of the world we live in with acceptance that there is no invisible hand. In this view, rational individual selfseeking behaviour is by no means the necessary and sufficient micro substrate for the desirability of social outcomes. Rather, behaviour must be social if desirable social outcomes are to be obtained.ă According to Keynes, for example, egotistical activity uncoordinated by the state may lead to inefficient outcomes. ă The price system aggregates rational individual actions but the aggregate is an unintended outcome as far as those individuals are concerned. There is no particular reason why unintended outcomes should necessarily be desirable and often they are not. ă Individuals take responsibility for maximising their own welfare, given what everyone else is doing, but society as a whole has to take responsibility for organising the aggregate outcome if undesirable aggregate outcomes are to be avoided: There is no design but our own ... the invisible hand is merely our own bleeding feet moving through pain and loss to an uncertain destination.Ť(Keynes, 1930, reprinted 1981: 471).

The verdict of this section, therefore, is that while reductionist approaches may be safely dismissed, embracing holism is no guarantee of getting it right. There is holism and there is holism. Methodological holism combined with the <i>deus ex machina </i>of an invisible hand mechanism can still sustain the inappropriate and unwarranted reductionist policy prescription of <i>laissezfaire</i>. The bulk of this paper has concerned the holism-reductionism dyad, at the cost of ignoring the idealism-materialism dyad. Nevertheless, as was argued earlier, holism can be paired with either materialism or idealism, and what Bunge, and I, advocate is, in my terminology, a materialist holism. The implication is that writers in the invisible hand tradition, such as Smith and Hayek, adopt an idealist from of holism. As far as Smith is concerned, the case has been made in Denis (1999) where the invisible hand is interpreted as literally the hand of an omniscient, omnipotent and benign deity. As far as Hayek is concerned, this is the topic of another paper, at the time of writing under consideration for publication by another journal. ă

The consequence of this analysis is that the association between methodological standpoint and policy prescription asserted in Bunge (2000) can be challenged. ă There is no simple association between ŚindividualismŠand laissez-faire, on the one hand, nor between ŚholismŠ and intervention or centralisation, on the other. ă On the contrary, <i>both </i>ŚindividualismŠ (reductionism) <i>and</i> ŚholismŠ (idealist holism) are associated with laissez-faire while ŚsystemismŠ (materialist holism) is associated with a range of policy prescriptions from mild to radical intervention.

4. Conclusion

In substance BungeŠs article is a significant contribution to our understanding of the methodology of economics Ű perhaps best embodied in his bold injunction to Śsee agency through WeberŠs microscope, and structure through MarxŠs telescopeŠ (154). ă The article, however, can be faulted in two respects. ă By compressing the two overlaying oppositions, between holism and reductionism and between idealism and materialism, into a single spectrum running from ŚindividualismŠ via ŚsystemismŠ to ŚholismŠ, he loses focus and clarity on the fundamental division between holistic and reductionistic world-views in economics and social science in general. ă This lack of clarity itself contributes to an over-simplification and misunderstanding of the relation between methodological premises and policy consequences. There do exist reductionist free marketeers of the kind Bunge describes: the pronouncements of Friedman, Lucas and Sir Keith fit this pattern. ă But to lump holists such as Smith and Hayek in with them, is both mistaken and allows us to ignore the fundamentally ideological role of invisible hand mechanisms in allowing economists to retain some approximation to efficiency as their default notion of how the capitalist economy actually works.